OWG Meeting 12-Mar-2014

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm

Remote Connection

ESNet: 8542553

(if problems, call phone in conference room: 757-269-6460)

Previous Meeting


  • Announcements
  • Major systems reports
    • Trigger
    • DAQ
    • Monitoring
    • Controls
  • E-log logbooks
  • Mantis Task Tracker


Attendees: David L. (chair), Hovanes E., Eugene C., Vardan G., Graham H., Bryan M., Mark I., Yi Q., Dave A., Sergey F., Simon T., Alex S., Sean D., Curtis M., Beni Z., Dmitry R., Sergey P.


Mark I. regarding DB re-organization:

  • Not too much interest in modifying existing DB schemes
  • Dropping efforts to reorganize for now


  • The FCAL North and South independent DAQ systems are now able to:
    • Use internal pulser for triggers
    • Use LED pulser
    • readout at ~10Hz level (full waveforms for all channels)
  • The BCAL North and South are also independent DAQ systems
    • The south 6 crates are connected to the TS
    • Able to read out all South crates
    • We are short on optical transceivers (need 12 more for BCAL)
      • unclear when order will arrive
      • need 2 transceivers per crate so they can have TS controlled and TI controller DAQ systems wired up simultaneously (although they can't be run simultaneously)
  • 8 TDs have been installed
  • Currently in process of checking timing calibration of trigger distribution system
  • Configuration file implementation nearly complete. Only a few things to finish, but other priorities have superseded completion of that project for the moment.
  • Hai has finished implementing code needed for the cosmic trigger in the flash 250.
  • SSPs are not yet connected


  • Sergey F.: Work has started on looking at overall organization of software directories and scripts
  • CODA & Linux VME drivers are currently being developed and released independently
    • CODA uses svn
    • Linux VME drivers use GIT
    • DAQ group willing to discuss tying the release of versions of the two together
  • CODA statistics are currently recorded to a file that Dmitry can (and does) read to store in DB


  • Multiple queries have come in regarding TTree objects in ROOTSpy
    • Hall-B is possibly interested
    • Hall-D fast-DAQ system used for solenoid.
      • currently writes values to ROOT file, but process must be stopped and ROOT file closed in order to read and access TTrees. Would like to do this in real-time


  • Dave. B is working on BCAL temperature monitoring
  • A Meeting was held to discuss the gas controls
  • Orlando finished the pulser control GUI
    • showed GUI to John Leckey and how to use it for FCAL
    • is currently incorporating John's feedback
  • BCAL voltage control is nearly complete
  • Nerses is now here
  • EEL setup for CDC is done, but likely will not be used as the CDC will be moved to Hall-D soon
  • * the chassis could be used for something else though
  • The Tagger magnet power supply controls software is now ready (handled by accelerator)
  • Yi is working on Solenoid PXI readout (modified for better readout)
  • Central Alarm System
    • Mark I.: Each group has been asked to provide a list of EPICS variables to alarm on
    • Hovanes: will use the FCAL darkroom as an example to start prototyping


  • e-logbooks:
    • There was some discussion on e-logbooks. Specifically, whether we should have multiple "logbooks" or maintain everything under a single logbook. Some the existing features of the e-logbook were shown to help aid the discussion.
    • The discussion was cut short by the chair after it was determined that a consensus would be unlikely.

Next Meeting

Next meeting is scheduled for March 26th